OpenStudy (anonymous):

I can't seem to solve this: In 8 years, Nicci will exceed Kahlan's age by 16 years. Last year, the sum of their ages was 20. How old is Nicci?

OpenStudy (anonymous):

Nikki's current age = N Kalhan's current age = K In 8 years' time : (N + 8) - ( K + 8) = 16 N - K = 16 ---(1) Last year : (N - 1) + (K - 1) = 20 N + K = 18 ---(2) Solve equations (1) and (2).

OpenStudy (anonymous):

Nicci is 19 years old

OpenStudy (anonymous):

I would like to point out that I think equation 2 is wrong. Shouldn't it be N+K=22?

OpenStudy (anonymous):

way off ! The answer is totally wrong, her's the correct one